To use the dual simplex method to solve the following LPP :

Minimise z=x₁+2x₂+3x₃

subject to

x₁-x₂+x₃>= 4

x₁+x₂+2x₃<= 8

x₁-x₃ >= 2

x₁,x₂, x₃>= 0


1
Expert's answer
2021-03-11T11:26:00-0500

Given inequality are-

x₁-x₂+x₃ "\ge" 4

x₁+x₂+2x₃"\le" 8

x₁-x₃ "\ge" 2

x₁,x₂, x₃ ">" 0


The Z-function is-

z=x₁+2x₂+3x₃


On solving The given inequality by eliminating "x_2" from eqn.(1) and (2) and simultaneously solving the equation with eqs.(3)

and we found the critical point i.e. (3.6,1.2,1.6)


The value of Z is equla to "3.6+2(1.2)+3(1.6)=3.6+2.4+4.8=10.8"


Hence The minimum value is 10.8